# A Pre-trained Sequential Recommendation Framework: Popularity Dynamics for Zero-shot Transfer

### 03 Jan 2024 v1

@@ Abstract.
Sequential recommenders are crucial to the success of online applications, e.g., e-commerce, video streaming, and social media. While model architectures continue to improve, for every new application domain, we still have to train a new model from scratch for high quality recommendations. On the other hand, pre-trained language and vision models have shown great success in zero-shot or few-shot adaptation to new application domains. Inspired by the success of pre-trained models in peer AI fields, we propose a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item representations by modeling item popularity dynamics. Through extensive experiments on five real-world datasets, we show that PrepRec, without any auxiliary information, can not only zero-shot transfer to a new domain, but achieve competitive performance compared to state-of-the-art sequential recommender models with only a fraction of the model size. In addition, with a simple post-hoc interpolation, PrepRec can improve the performance of existing sequential recommenders on average by 13.8% in Recall@10 and 29.5% in NDCG@10. We provide an anonymized implementation of PrepRec at https://anonymous.4open.science/r/PrepRec--2F60/.
@@ 1. Introduction
Modeling sequential user behavior is critical to the success of online applications such as e-commerce, video streaming, and social media. Despite essential innovations for tackling the sequential recommendation task (Rendle et al., 2010; Hidasi et al., 2015; Tan et al., 2016; Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020), these models have to be trained from scratch for each application domain. because they learn representations for each item (Sun et al., 2019; Kang and McAuley, 2018) in the domain. Further, to maintain high performance, we need to retrain the models when there is a large influx of new users and items. Since the models learn representations for each item, they cannot generalize to new domains. Thus, in this paper, we ask: can we build a pre-trained sequential recommender system capable of cross-domain and cross-application zero-shot transfer without any auxiliary information? (e.g., using the model trained for online shopping in the US to predict the next movie that a user in India will watch).
@@
At first glance, developing pre-trained sequential recommenders for zero-shot cross-domain inference seems impossible. While we see pre-trained language (Devlin et al., 2018b; Brown et al., 2020; OpenAI, 2023; Raffel et al., 2020; Liu et al., 2019) and vision models (Dosovitskiy et al., 2020; He et al., 2016; Radford et al., 2021) show excellent generalizability across datasets and applications, being able to achieve state-of-the-art performance by just a few fine-tuning steps (Devlin et al., 2018b; Liu et al., 2019) or even without any training (OpenAI, 2023; Brown et al., 2020) (i.e., zero-shot transfer), there are are essential differences. The representations learned by the pre-trained language model seem universal since the training domain and the application domain (e.g., text prediction and generation) share the same language and vocabulary, supporting the effective reuse of the word representations. However, in the cross-domain recommendation problem, the items are distinct across domains in sequential recommendation datasets (e.g., grocery items vs movies). Therefore, forming such generalizable correspondence is nearly impossible if we learn specific representations for each item within each domain. Recent work explores pre-trained models for sequential recommendation (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021) within the same application (e.g., online retail). However, they assume access to auxiliary information of items (e.g., item description), which is application-dependent and is not always available in practice. Therefore, these models cannot learn universal representations of items; instead, they bypass the representation learning problem by using additional item side information. Thus, while achieving success within a domain, these recent models cannot generalize to new domains without auxiliary information. This paper tackles the challenge of zero-shot, cross-domain sequential recommendations without any auxiliary information.
@@
Present Work: Our crucial insight is to learn the universal item representations based on the item popularity dynamics and use these representations for cross-domain, zero-shot prediction. We know that the marginal distribution of user and item activities is universal across datasets, supported by prior work in network science (Barabási and Albert, 1999; Barabasi, 2005) and by experiments in recommender systems (Salganik et al., 2006). In addition, recent work in recommender systems suggests that the popularity dynamics of items are also crucial for predicting users’ behaviors (Ji et al., 2020). Therefore, in this paper, we explicitly model the popularity dynamics of items and propose a novel pre-trained sequential recommendation framework: PrepRec. We learn to represent items based on their popularity dynamics, instead of their explicit item IDs. We encode the time interval between two consecutive interactions via relative-time encoding and the position of each interaction in the sequence by positional encoding. In contrast, prior work in sequential recommendation uses logical indices. Using physical time ensures that the predictions are not anti-causal, i.e., using the future interactions to predict the present. We learn the representations within a transformer architecture coupled with optimizations. The learned item popularity representations, time-interval and positional encoding are all universal, making it possible to build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information. Our key contributions are as follows:
Universal item representations::
@@
We are the first to learn universal item representations for sequential recommendation. In contrast, prior research learns item representations for each item ID or through item auxiliary information. We learn universal item representations by exploiting item popularity dynamics. We learn two temporal representations using a transformer architecture with optimizations at any time $t$ for each item’s popularity: at a coarse and fine-grained level. We represent items’ popularity dynamics (i.e., representing popularity changes) by concatenating representations over a fixed time interval. Item dynamics are inferrable from the user-item interaction data, and thus, the learned item representations are transferable across domains and applications. These item representations make possible pre-trained sequential recommender systems capable of cross-domain and cross-application transfer without any auxiliary information.
@@
We propose a new challenging setting for pre-trained sequential recommender systems: zero-shot transfer without any auxiliary information. In contrast, previous works in sequential recommender systems capable of cross-domain zero-shot rely heavily on application-dependent auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021). To the best of our knowledge, we are the first to tackle this challenging setting in sequential recommendation.
With extensive experiments, we empirically show that PrepRec has excellent generalizability across domains and applications. Remarkably, had we trained a state of the model from scratch for the target domain, instead of zero-shot transfer using PrepRec , the performance gain would have been only (on average) 0.2%. In addition, we show that with a post-hoc interpolation, PrepRec can outperform the state-of-the-art sequential recommender system on average by 13.8% in Recall@10 and 29.5% in NDCG@10. With this work, it becomes clear that item popularity dynamics explain much of the prediction ability of the modern neural recommender systems, shedding light on what neural models are learning from data and opening a new direction for improvement in the future.
@@ 2. Related Work
Sequential Recommendation: Sequential recommenders model user behavior as a sequence of interactions, and aim to predict the next item that a user will interact. Early sequential recommenders adopt Markov chains and model item-item transition probability (Rendle et al., 2010; Shani et al., 2005). Later, with the advance of neural networks, some basic neural architectures have been explored for sequential recommendation, such as convolutional neural networks (Tang and Wang, 2018; Tuan and Phuong, 2017). Recurrent neural networks (Hidasi et al., 2016) and its variants, such as GRU4Rec (Hidasi et al., 2015) and GRU4Rec+ (Tan et al., 2016), have also been widely used for sequential recommendation due to the sequential nature of RNN-based architectures.
@@
Recently, with the success of attention mechanisms in modeling sequential data, e.g., text sequences, some works have explored using attention mechanisms for sequential recommendation (Li et al., 2017; Liu et al., 2018). Attention-based models, e.g., Transformers (Vaswani et al., 2017) and BERT (Devlin et al., 2018a) show state-of-the-art performance in many NLP tasks. Inspired by the success of Transformers and BERT, some works have explored using attention-based models for sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020). However, these works only focus on the regular sequential recommendation and do not consider the zero-shot sequential recommendation. Some works take the timestamps of each interaction into consideration. (Li et al., 2020) looks at the time interval between two consecutive interactions and proposes a time-aware attention mechanism for sequential recommendation. Others (Lv et al., 2019; Ying et al., 2018; Tan et al., 2021) separate interaction sequences and categorize them to show the long-term and short-term interests of users. Some work on temporal recommendation model the change in users’ preferences and item popularities (Koren, 2009; Wu et al., 2017; Zhang et al., 2014). However, these works do not consider the dynamics of item popularity statistics and are also not capable of zero-shot sequential recommendation.
Symbol Description
$P(a^{t}_{j})$ popularity percentile of item $v_{j}$ at time $t$ over coarser time period
$P(b^{t}_{j})$ popularity percentile of item $v_{j}$ at time $t$ over finer time period
$\mathcal{P}^{t}_{j}$ popularity dynamics of item $v_{j}$ over a coarser time period up to time $t$
$\mathcal{H}^{t}_{j}$ popularity dynamics of item $v_{j}$ over a finer time period up to time $t$
$\mathbf{e}^{t}_{j}$ popularity dynamics embedding of item $v_{j}$ at time $t$
${\bm{T}}$ relative time interval encoding matrix
${\bm{P}}$ positional encoding matrix
$\mathbf{q}_{u}$ user embedding, computed from § 4.1.4
$h$ number of attention heads
$d$ embedding dimension
Table 1. Notation Table
@@
Transfer Learning in Recommendation: Cross-domain transfer learning has been explored in recommender systems. Some works in the cross-domain recommendation literature focus on leveraging the relatively information-rich domain to improve the recommendation performance on the data-sparse domain (Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017). However, most of these works assume user or item overlap (Zhu et al., 2021; Zhao et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017) for effective knowledge transfer. Other cross-domain literature focuses on the cold-start problem. There have been works proposed to tackle the cold-start user case (Lu et al., 2020; Zhu et al., 2021; Yuan et al., 2020; Feng et al., 2021) or the cold-start item case (Volkovs et al., 2017; Wei et al., 2021; Du et al., 2020; Liu et al., 2021). However, these works assume auxiliary information is available and can only handle one-sided cold-start problems (i.e., either cold-start users or cold-start items). Some other past works explore meta-learning on the cold-start problem and can handle both cold-start users and cold-start items (Lee et al., 2019; Dong et al., 2020) with application-dependent side information and fine-tuning on the target domain. Note that in this work, we are not trying to address the cold-start problem. Instead, we focus on the zero-shot transfer problem, where no auxiliary information is available in both domains.
@@
Recently, pre-trained recommenders capable of zero-shot transfer have caught some attention in the community. (Ding et al., 2021; Hou et al., 2022) propose pre-trained models for sequential recommendations. However, they rely heavily on dataset-dependent auxiliary information (i.e., description text of items). (Wang et al., 2023) investigates the joint and marginal activity distribution of users and items, but are not suitable for the sequential recommendation task.
@@
To summarize, prior works on sequential recommendation focus on learning high-quality representations for each item in the training set and are not generalizable across domains. Pre-trained sequential recommenders are capable of zero-shot recommendation to a new domain but rely heavily on auxiliary information of items, which is not always available in practice. In this paper, we propose a novel pre-trained sequential recommendation framework based on item popularity dynamics that are universal across datasets, and it is capable of zero-shot transfer without any auxiliary information.
@@ 3. Problem Definition
In this section, we formally define the research problems this paper addresses (i.e., regular sequential recommendation and zero-shot sequential recommendation) and introduce the notations used in this paper.
In this paper, we consider building the zero-shot sequential recommender system with only implicit feedback, i.e., no auxiliary information. In sequential recommendation, denote ${\bm{M}}$ as the implicit feedback matrix, $\mathcal{U}$ = $\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ as the set of users, $\mathcal{V}$ = $\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ as the set of items. The goal of sequential recommendation is to learn a scoring function, that predicts the next item $v_{u,t+1}$ given a user $u$ ’s interaction history $\{v_{u,1},v_{u,2},...,v_{u,t}\}$. Note that in this paper, since we model time explicitly, we assume access to the timestamp of each interaction, including the next item interaction. We argue that this is a reasonable assumption since the timestamp of the next interaction is always available in practice. For example, if Alice logs in to Netflix, Netflix will always know when Alice logs in and can predict the next movie for Alice. Formally, we define the scoring function as $\mathcal{F}(v^{t}|\mathcal{S}_{u},{\bm{M}})$, where $t$ is the time of the prediction.
Zero-shot Sequential Recommendation: In zero-shot sequential recommendation, given an interaction matrix ${\bm{M}}^{\prime}$ over $\mathcal{U}^{\prime}$ and $\mathcal{V^{\prime}}$, the goal is to produce a scoring function $\mathcal{F}^{\prime}$ without training on ${\bm{M}}^{\prime}$. In other words, the scoring function $\mathcal{F}^{\prime}$ has to be trained on a different interaction matrix ${\bm{M}}$ and then applied to ${\bm{M}}^{\prime}$ directly to predict the next item that user $u$ will interact with. In this paper, we study the zero-shot sequential recommendation that not only ${\bm{M}}$ and ${\bm{M}}^{\prime}$ are from different domains, but also both $\mathcal{U}\cap\mathcal{U}^{\prime}=\emptyset$ and $\mathcal{V}\cap\mathcal{V}^{\prime}=\emptyset$. In addition, we assume no auxiliary information is available in both domains.
Compared to previous works on zero-shot sequential recommendation, which assume access to side information, we pose a more challenging problem setting. We argue that this experimental setting is more universal and practical, and requires the model to learn the truly universal properties of the data.

### 10 Apr 2024 v2

# Pre-trained Sequential Recommendation Framework via Popularity Dynamics

@@ Abstract.
Sequential recommenders models sequential users’ behavior and are the fundamental building blocks of many online applications, e.g., e-commerce, video streaming, and social media. While effective, they learn item embeddings one-on-one mapped to explicit item IDs, which are unique in different domains, making them hard to generalize to new recommendation domains, whereas pre-trained models have shown great success in generalizing to new application domains in peer AI fields. Pioneer works on pre-trained sequential recommenders rely on application-dependent auxiliary information and can only generalize to the related domains within the same application, e.g., online retail. In this work, we aim to set a baseline for pre-trianed sequential recommenders by tackling the most challenging task: zero-shot cross-domain and cross-application sequential recommendation without auxiliary information. We present a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item and sequence representations by modeling item dynamics and propose a popularity dynamics-aware transformer architecture. Through extensive experiments on five real-world datasets, we show that PrepRec, without any auxiliary information, can zero-shot adapt to new application domains and achieve competitive performance compared to state-of-the-art sequential recommender models. In addition, with a simple post-hoc interpolation, PrepRec can improve the performance of existing sequential recommenders on average by 12.4% in Recall@10 and 22% in NDCG@10. We provide an anonymized implementation of PrepRec at https://anonymous.4open.science/r/PrepRec--2F60/.
@@ 1. Introduction
Modeling sequential user behavior is critical to the success of online applications such as e-commerce, video streaming, and social media. Despite essential innovations for tackling the sequential recommendation task (Rendle et al., 2010; Hidasi et al., 2015; Tan et al., 2016; Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020), these models have to be trained from scratch for each application domain because they learn representations for each item (Sun et al., 2019; Kang and McAuley, 2018) in the domain. Further, to maintain high performance, we need to retrain the models when there is a large influx of new users and items. Since the models learn representations for each item, they cannot generalize to new domains. Thus, in this paper, we ask: can we build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information? (e.g., using the model trained for online shopping in the US to predict the next movie that a user in India will watch).
@@
At first glance, developing pre-trained sequential recommenders for cross-domain inference seems impossible. While we see pre-trained language (Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023; Raffel et al., 2020; Liu et al., 2019) and vision models (Dosovitskiy et al., 2020; He et al., 2016; Radford et al., 2021) show excellent generalizability across datasets and applications, being able to achieve state-of-the-art performance by just a few fine-tuning steps (Devlin et al., 2018; Liu et al., 2019) or even without any training (OpenAI, 2023; Brown et al., 2020) (i.e., zero-shot transfer), there are essential differences. The representations learned by the pre-trained language model seem universal since the training domain and the application domain (e.g., text prediction and generation) share the same language and vocabulary, supporting the effective reuse of the word representations. However, in the cross-domain recommendation problem, the items are distinct across domains in sequential recommendation datasets (e.g., grocery items vs movies). Therefore, forming such generalizable correspondence is nearly impossible if we learn specific representations for each item within each domain. Recent work explores pre-trained models for sequential recommendation (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021) within the same application (e.g., online retail). However, they assume access to metadata of items (e.g., item description), which is application-dependent and is often not generalizable to other applications. These models cannot learn universal representations of items; instead, they bypass the representation learning problem by using additional item-side information. This paper tackles the more challenging task of zero-shot, cross-domain and cross-application sequential recommendations without any auxiliary information.
@@
Present Work: Our crucial insight is to learn the universal item representations based on the item popularity dynamics and use these representations for cross-domain and cross-application zero-shot prediction. We know that the marginal distribution of user and item activities is universal and heavy tailed across datasets, supported by prior work in network science (Barabási and Albert, 1999; Barabasi, 2005) and by experiments in recommender systems (Salganik et al., 2006). In addition, recent work in recommender systems suggests that the popularity dynamics of items are also crucial for predicting users’ behaviors (Ji et al., 2020). Therefore, in this paper, we explicitly model the popularity dynamics of items and propose a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item representations based on their popularity dynamics, instead of their explicit item IDs or auxiliary information. We encode the relative time interval between two consecutive interactions via relative-time encoding and the position of each interaction in the sequence by positional encoding. Using physical time ensures that the predictions are not anti-causal, i.e., using the future interactions to predict the present. We propose a popularity dynamics-aware transformer architecture for learning universal sequence representations. We show that it is possible to build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information. Our key contributions are as follows:
Universal item and sequence representations::
@@
We are the first to learn universal item and sequence representations for sequential recommendation without any auxiliary information by exploiting item popularity dynamics. In contrast, prior research learns item representations for each item ID or through item auxiliary information. We learn universal item representations by modeling item popularity dynamics of two temporal resolutions: coarse and fine-grained. We learn universal sequence representations using a carefully designed popularity dynamics-aware transformer architecture. These universal item and sequence representations make possible pre-trained sequential recommender systems capable of cross-domain and cross-application transfer without any auxiliary information.
@@
We propose a new challenging setting for pre-trained sequential recommender systems: zero-shot cross-domain and cross-application transfer without any auxiliary information. In contrast, previous pre-trained sequential recommenders requires overlapping users (Yuan et al., 2020), application-dependent auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021; Hou et al., 2023; Wang et al., 2023b), and are few-shot adapted to related domains within the same application (Hou et al., 2022, 2023; Wang et al., 2023b). Assuming no user or item metadata posts a more challenging but crucial problem as we want to set a baseline for the performance of pre-trained sequential recommenders. Using metadata can only make the problem easier and improve the performance.
With extensive experiments, we empirically show that PrepRec has excellent generalizability across domains and applications. Remarkably, had we trained a state-of-the-art model from scratch for the target domain, instead of zero-shot transfer using PrepRec, the maximum performance gain would have been only 4%. In addition, we show that with a post-hoc interpolation, PrepRec can outperform the state-of-the-art sequential recommender system on average by 12.4% in Recall@10 and 22% in NDCG@10. We attribute the improvements to the performance gains over long-tail items, which we show in the qualitative analysis. With this work, we set a baseline for pre-trained sequential recommenders and show that it is possible to build a pre-trained sequential recommender system under the most challenging setting: cross-domain and cross-application transfer without any auxiliary information.
@@ 2. Related Work
Sequential Recommendation: Sequential recommenders model user behavior as a sequence of interactions, and aim to predict the next item that a user will interact with. Early sequential recommenders adopt Markov chains and model item-item transition probability (Rendle et al., 2010; Shani et al., 2005). Later, with the advance of neural networks, some basic neural architectures have been explored for sequential recommenders, such as convolutional neural networks (Tang and Wang, 2018; Tuan and Phuong, 2017). Recurrent neural networks (Hidasi et al., 2016) and its variants, such as GRU4Rec (Hidasi et al., 2015) and GRU4Rec+ (Tan et al., 2016), have also been widely used for sequential recommendation due to the sequential nature of RNN-based architectures.
@@
With the success of Transformer (Vaswani et al., 2017) in modeling sequential data, e.g., text sequences, some works adopt the transformer architecture for sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020). Additionally, (Li et al., 2020) considers the timestamps of each interaction and looks at the time interval between two consecutive interactions and proposes a time-aware attention mechanism. Some works (Lv et al., 2019; Ying et al., 2018; Tan et al., 2021) separate interaction sequences and categorize them to show the long-term and short-term interests of users. Temporal sequential recommenders (Koren, 2009; Wu et al., 2017; Zhang et al., 2014) models the change in users’ preferences. These works, while achieving state-of-the-art performance, only focus on the regular sequential recommendation and cannot transfer to other domains.
@@
Cross-domain Recommendation: Cross-domain recommendation literature leverages the information-rich domain to improve the recommendation performance on the data-sparse domain (Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017). However, most of these works assume user or item overlap (Zhu et al., 2021; Zhao et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017) for effective knowledge transfer. Other cross-domain literature focuses on the cold-start problem (Lu et al., 2020; Zhu et al., 2021; Feng et al., 2021; Volkovs et al., 2017; Wei et al., 2021; Du et al., 2020; Liu et al., 2021; Lee et al., 2019; Dong et al., 2020). Note that we are not trying to address the cold-start problem. Instead, we focus on building a pre-trained sequential recommendation framework that can zero-shot transfer to a new domain without any auxiliary information.
@@
Pre-trained Sequential Recommenders: Recently, with the success of pre-trained language models(Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023), pre-trained recommenders have caught the attention of the community. ZESRec (Ding et al., 2021) is capable of zero-shot sequential recommendations. However, it only works for closely related domains and relies on application-dependent auxiliary information (i.e., description text of items). PeterRec (Yuan et al., 2020) requires overlapping users in both domains. On the other hand, finetuning-based models, e.g., MISSRec (Wang et al., 2023b), UnisRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023), are not designed for zero-shot sequential recommendation and works within the same application (e-commerce) since they rely on application-dependent auxiliary information.  (Wang et al., 2023a) investigates the joint and marginal activity distribution of users and items, but are not suitable for the sequential recommendation task.
@@
To summarize, prior works on sequential recommendation focus on learning high-quality representations for each item in the training set and are not generalizable across domains. Pre-trained sequential recommenders are evaluated on closely related domains and platforms and rely heavily on application-dependent auxiliary information of items.
@@ 3. Problem Definition
In this section, we formally define the research problems this paper addresses (i.e., regular sequential recommendation and zero-shot sequential recommendation) and introduce our notations.
In sequential recommendation, denote ${\bm{M}}$ as the implicit feedback matrix, $\mathcal{U}$ = $\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ as the set of users, $\mathcal{V}$ = $\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ as the set of items. The goal of sequential recommendation is to learn a scoring function, that predicts the next item $v_{u,t}$ given a user $u$ ’s interaction history $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,t-1}\}$. Note that in this paper, since we model time explicitly, we assume access to the timestamp of each interaction, including the next item interaction. We argue that this is a reasonable assumption since the timestamp of the next interaction is always available in practice. For example, if Alice logs in to Netflix, Netflix will always know when Alice logs in and can predict the next movie for Alice. Formally, we define the scoring function as $\mathcal{F}(v^{t}|\mathcal{S}_{u},{\bm{M}})$, where $t$ is the time of the prediction.
Zero-shot Sequential Recommendation: Given two domains ${\bm{M}}$ and ${\bm{M}}^{\prime}$ over $\mathcal{U},\mathcal{V}$ and $\mathcal{U}^{\prime},\mathcal{V^{\prime}}$ respectively, we study the zero-shot recommendation problem in the scenario where the domains are different (${\bm{M}}\cap{\bm{M}}^{\prime}=\varnothing$), users are disjoint ($\mathcal{U}\cap\mathcal{U^{\prime}}=\varnothing$), and item sets are unique ($\mathcal{V}\cap\mathcal{V^{\prime}}=\varnothing$). The goal is to produce a scoring function $\mathcal{F}^{\prime}$ without training on ${\bm{M}}^{\prime}$ directly. In other words, the scoring function $\mathcal{F}^{\prime}$ has to be trained on a different interaction matrix ${\bm{M}}$. Furthermore, we assume there is no metadata associated with users or items, which makes the problem particularly challenging but crucial to study. We want to set a baseline for pre-trained sequential recommenders, using metadata can only improve and simply the problem.